# A Unified Voxel Diffusion Module for Point Cloud 3D Object Detection

### v1

@@ Introduction
VDM-OD-Mamba (Ours) 81.3/80.9 73.0/72.6 86.1/82.1 78.9/75.0 80.7/79.7 77.8/76.9 76.6/74.8
Table 2: Comparison with prior methods (CNNs, Transformers, and SSMs) on the Waymo Open Dataset using 100% training data. Metrics are reported as mAP/mAPH (%) $\uparrow$ for overall results and AP/APH (%) $\uparrow$ per category. All models are trained under the single-frame setting. “–” indicates unavailable results. VDM-OD-Mamba denotes a variant where the Only-Diffusion structure is used within the VDM framework. The best-performing results are highlighted in bold.
Fine-Grained Feature Aggregation. To capture more detailed spatial information, we further divide the voxel grid into finer partitions, as illustrated in Figure 4. Leveraging the strong spatial perception capabilities of the Submanifold 3D Convolution (SubM3D) and the Sparse Residual Block (SRB) within the VDM, we extract fine-grained voxel features from these subdivided regions. In addition, we adopt a sparse convolution with a stride of 2 to serve dual purposes: first, it reduces the resolution of the feature map to half, which effectively controls computational cost by mitigating the overhead introduced by voxel diffusion; second, it aggregates local fine-grained features, thereby enriching the spatial context available for downstream modules. This design ensures that the model benefits from both detailed spatial representations and efficient computation during object detection.
The function of VDM is to diffuse the number of sparse voxels and to aggregate fine-grained features. It can be summarized by the following formulas:
$\mathit{FG}^{\prime}=\mathrm{Voxel\_Diffusion}(\mathit{FG},\mathit{spconv})$ (6)
$\mathit{F}^{\prime}=\mathrm{Aggregation}(\mathit{F},\mathit{subm},\mathit{spconv})$ (7)
where $\mathit{F}$ denotes the initial voxel feature map; $\mathit{F}^{\prime}$ represents the feature map after local voxel aggregation, with width and height reduced to one-fourth of the original size; $\mathit{subm}$ denotes submanifold sparse convolution, and $\mathit{spconv}$ denotes sparse convolution. $\mathit{FG}^{\prime}$ indicates the number of foreground voxels after VDM processing, and $\mathit{FG}$ is the number of foreground voxels in the initial feature map.
Integration with SSMs and Transformer Modules. The proposed VDM module can be integrated with both Mamba-based and Transformer-based models. By effectively diffusing sparse voxels and aggregating fine-grained point cloud voxel features, VDM further enhances the detection accuracy of the overall model. For the 3D feature maps processed by VDM, we first serialize them into one-dimensional sequences along the $x$ and $y$ directions. As the sequence length increases, Mamba-based networks are able to adopt larger grouping sizes due to their lower computational complexity compared to Transformer-based counterparts. Given the strong performance of LION (Liu et al. 2024) and DSVT (Wang et al. 2023) in the field of 3D object detection, we integrate the VDM module into both networks and conduct experiments to validate its effectiveness.
Experiments
Datasets and Evaluation Metrics
Waymo Open Dataset. The Waymo Open Dataset (WOD) (Sun et al. 2020) includes 1,150 scenes, divided into 798 for training, 202 for validation, and 150 for testing. Each scene covers a perception range of $150\,\text{m}\times 150\,\text{m}$. For evaluation, WOD adopts 3D mean Average Precision (mAP) and mAP weighted by heading accuracy (mAPH). Each metric includes two difficulty levels: L1 for objects with more than five points, and L2 for bounding boxes containing one to five points.
nuScenes Dataset. The nuScenes dataset (Caesar et al. 2020) contains 1,000 scenes, with 750 for training, 150 for validation, and 150 for testing. It uses mean Average Precision (mAP) and nuScenes Detection Score (NDS) as evaluation metrics. The dataset covers 10 annotated object classes, representing most common categories in traffic scenarios.
Argoverse 2 Dataset. The Argoverse 2 dataset (Chang et al. 2019) comprises 1,000 sequences, with 700 for training, 150 for validation, and 150 for testing. The perception range extends up to 200 meters. It uses mean Average Precision (mAP) as the evaluation metric.
ONCE Dataset. The ONCE dataset (Mao et al. 2021) consists of 5,000, 3,000, and 8,000 frames for training, validation, and testing respectively. It defines three detection classes: Vehicle, Pedestrian, and Cyclist. The final performance is evaluated using mean Average Precision (mAP) across all three classes.
@@
CenterPoint 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
VoxelNeXt 66.7 60.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
Uni3DETR 68.5 61.7
Transfusion-LiDAR 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
DSVT 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.9 71.0
HEDNet 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
SAFDNet 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
LION 72.1 68.0 87.9 64.9 77.6 44.4 28.5 89.6 75.6 59.4 80.8 71.6
UniMamba 72.6 68.5 88.7 64.7 79.7 47.9 28.7 89.7 74.6 59.1 79.5 72.3
@@
TransFusion-LiDAR 70.2 65.5 86.2 56.7 66.3 58.8 28.2 86.1 68.3 44.2 82.0 78.2
DSVT 72.7 68.4 86.8 58.4 67.3 63.1 37.1 88.0 73.0 47.2 84.9 78.4
HEDNet 72.0 67.7 87.1 56.5 70.4 63.5 33.6 87.9 70.4 44.8 85.1 78.1
VPF 72.7 67.0 85.8 55.1 63.5 62.1 33.3 87.6 72.5 48.6 82.9 78.2
SAFDNet 72.3 68.3 87.3 57.3 68.0 63.7 37.3 89.0 71.1 44.8 84.9 79.5
LION 73.9 69.8 87.2 61.1 68.9 65.0 36.3 90.0 74.0 49.2 87.3 79.5
@@
Table 3: Performances on the nuScenes validation and test set. ‘T.L.’, ‘C.V.’, ‘Ped.’, ‘M.T.’, ‘T.C.’, and ’B.R.’ are short for trailer, construction vehicle, pedestrian, motor, traffic cone, and barrier, respectively. VDM-OD-Mamba denotes a variant where the Only-Diffusion structure is used within the VDM framework.
@@
CenterPoint 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
HEDNet 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
VoxelNeXt 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
FSDV2 37.6 77.0 47.6 70.5 43.6 41.5 53.9 58.5 56.8 39.0 60.7 49.4 28.4 41.9 30.2 44.9 33.4 16.6 7.3 32.5 45.9 24.0 1.0 12.6 17.1 26.3 17.2
SAFDNet 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
LION 41.5 75.1 43.6 73.9 53.9 45.1 66.4 74.7 61.3 48.7 65.1 56.2 21.7 42.7 25.3 58.4 28.9 23.6 8.3 49.5 47.3 19.0 0.0 31.4 8.7 37.6 11.8
UniMamba 42.0 78.9 47.9 74.3 51.8 46.8 67.8 76.9 55.8 51.7 62.8 52.2 30.2 44.6 24.6 28.1 59.4 32.2 23.2 6.7 41.5 48.5 0.0 26.4 8.1 36.4 13.7
@@
Table 4: Comparison with prior methods on the Argoverse2 validation set. ‘Vehicle’, ‘C-Barrel’, ‘MPC-Sign’, ‘A-Bus’, ‘C-Cone’, ‘V-Trailer’, ‘MBT’, ‘W-Device’, and ‘W-Rider’ are abbreviations for regular vehicle, construction barrel, mobile pedestrian crossing sign, articulated bus, construction cone, vehicular trailer, message board trailer, wheeled device, and wheeled rider, respectively. VDM and VDM-OD refer to the proposed VDM-Mamba and VDM-OD-Mamba variants.
Implementation Details
Network Architecture. We evaluate the proposed VDM module by integrating it with both the SSMs-based LION (Liu et al. 2024) and the Transformer-based DSVT (Wang et al. 2023) models. We use voxel sizes of $(0.075\,\text{m},0.075\,\text{m},0.25\,\text{m})$, $(0.08\,\text{m},0.08\,\text{m},0.1875\,\text{m})$, $(0.1\,\text{m},0.1\,\text{m},0.25\,\text{m})$, and $(0.1\,\text{m},0.1\,\text{m},0.25\,\text{m})$ for the nuScenes, Waymo, Argoverse 2, and ONCE datasets respectively. The number of voxel feature channels is set to 64.
In the VDM module, the SubM3D (including SRB) layers use a kernel size of 3. All SparseConv3D layers adopt a kernel size of 3, stride of 2, and padding of 1 to downsample the feature map resolution. The channel configurations within VDM are set as $(64,32)$, $(32,64)$, and $(64,128)$.
For LION and DSVT, we follow the default parameter settings from their original papers without modification. Additionally, we conduct an only-diffusion experiment, in which the voxel size along the $x$ and $y$ axes is scaled by a factor of 4, while the $z$ axis remains unchanged.
Training Process. We train all models using 8 NVIDIA A40 GPUs. On the Waymo Open Dataset, the VDM-SSMs configuration is trained for 24 epochs with a batch size of 16, while the VDM-DSVT configuration uses 24 epochs with a batch size of 8. For the remaining three datasets, we adopt the same training strategy as that used in the LION (Liu et al. 2024) baseline.
Main Results
@@
Results on WOD. As shown in Table 2, we incorporate the VDM module into both Mamba-based and Transformer-based models. Among them, VDM-Mamba achieves a state-of-the-art (SOTA) result of 74.8 mAPH (L2), outperforming the corresponding baseline LION model (without VDM) by +0.7 mAPH, and also surpassing the recent UniMamba (Jin et al. 2025).
In addition, VDM-DSVT also demonstrates performance gains, achieving a +1.2 mAPH (L2) improvement over the original DSVT, which validates the general applicability of the VDM module to both SSM-based and Transformer-based detectors.
When only the diffusion mechanism is retained in VDM (i.e., VDM-OD-Mamba), the model still achieves the best results on the Vehicle and Cyclist categories, highlighting the importance of the diffusion component in the overall design. Moreover, VDM-Mamba achieves the best performance in the Pedestrian category, suggesting that the fine-grained aggregation ability introduced by VDM is particularly beneficial for small object detection.
@@
PointRCNN 52.1 74.5 40.9 16.8 4.3 6.2 2.4 0.9 29.8 46.0 20.9 5.5 28.7
PointPillars 68.6 80.9 62.1 47.0 17.6 19.7 15.2 10.2 46.8 58.3 40.3 25.9 44.3
SECOND 71.2 84.0 63.0 47.3 26.4 29.3 24.1 18.1 58.0 70.0 52.4 34.6 51.9
PV-RCNN 77.8 89.4 72.6 58.6 23.5 25.6 22.8 17.3 59.4 71.7 52.6 36.2 53.6
CenterPoint 66.8 80.1 59.6 43.4 49.9 56.2 42.6 26.3 63.5 74.3 57.9 41.5 60.1
PointPainting 66.2 80.3 59.8 42.3 44.8 52.6 36.6 22.5 62.3 73.6 57.2 40.4 57.8
LION 78.2 89.1 72.6 57.5 53.2 62.4 44.0 24.5 68.5 79.2 63.2 47.1 66.6
VDM-mamba 78.6 89.1 74.3 59.3 54.2 63.3 43.2 26.1 69.9 79.4 64.9 49.8 67.6
Table 5: Comparison with previous methods on ONCE validation set.
@@
Results on nuScenes. To further verify the effectiveness of the proposed VDM module, we conduct experiments on both the nuScenes validation and test sets. As shown in Table 3, VDM-mamba achieves 68.1 mAP and 72.5 NDS on the validation set, and 70.0 mAP and 73.7 NDS on the test set.
Although VDM-mamba slightly underperforms UniMamba in terms of overall detection accuracy, it achieves the best results in the pedestrian and bicycle categories on the nuScenes validation set. These categories are particularly important for autonomous driving applications, demonstrating that our proposed method offers clear advantages in detecting small and vulnerable objects. Furthermore, VDM-OD-Mamba achieves a state-of-the-art 72.9 NDS, suggesting that the diffusion-only structure plays a critical role in improving detection performance.
@@
Results on Argoverse V2. As shown in Table 4, our proposed models, VDM-Mamba and VDM-OD-Mamba, achieve state-of-the-art detection performance on the Argoverse 2 validation set, reaching 42.3 mAP and 42.6 mAP, respectively. Compared with the strong baseline LION (Liu et al. 2024), VDM-Mamba yields a notable improvement of 0.8 mAP. Notably, VDM-OD-Mamba, which retains only the diffusion structure within the VDM module, still surpasses all previous approaches, highlighting the effectiveness of the diffusion mechanism. These results further demonstrate the strong representational capacity and scalability of the VDM module in large-scale point cloud scenarios.
Results on ONCE. VDM-Mamba also demonstrates superior performance on the ONCE dataset. It achieves detection accuracies of 78.6 mAP, 54.2 mAP, and 69.9 mAP on the Vehicle, Pedestrian, and Cyclist categories, respectively. The overall detection accuracy across all categories reaches 67.6 mAP, representing a +1.0 mAP improvement over the Mamba-based LION detector. These results establish VDM-Mamba as the new state-of-the-art (SOTA) on the ONCE benchmark.
Ablation Study
Diffusion Waymo(3D AP/APH)L2
Vehicle Pedestrian Cyclist
No(LION) 68.8/68.3 77.3/72.6 74.7/73.8
Yes(VDM-OD) 70.3/69.9 78.2/73.8 76.0/75.0
Table 6: Detection results of the voxel diffusion module on the Waymo validation set (20% data). ”No” indicates that the original voxels are not diffused, and the detection results are entirely determined by LION(Liu et al. 2024) ”Yes” indicates that we only adopt the diffusion module from VDM, i.e., using sparse 3D spconv with a kernel size of 3 and a stride of 1.
Ablation Study on Voxel Diffusion. To assess the contribution of voxel diffusion to detection performance, we conduct ablation experiments on the Waymo validation set using a fixed voxel size of [0.32, 0.32, 0.1875]. We evaluate the model both with and without voxel diffusion. As shown in Table 6, introducing voxel diffusion improves the VDM-OD model by 1.3 mAPH (L2), underscoring its critical role in sequential detection frameworks. Moreover, on the nuScenes (Table 3) and Argoverse 2 (Table 4) datasets, we observe that under the Only-diffusion setting, the model still achieves higher detection accuracy, further validating the significant impact of voxel diffusion.
@@
Table 7: Ablation study on the ONCE and Waymo (20% data) dataset for fine-grained feature aggregation. Base denotes the detection results obtained by combining the VDM and Mamba network without fine voxel partitioning. Fine represents the results after applying fine-grained voxel partitioning.

### v3

@@ III-B Overall Architecture
VDM-OD (Ours) 81.3/80.9 73.0/72.6 86.1/82.1 78.9/75.0 80.7/79.7 77.8/76.9 76.6/74.8
TABLE II: Comparison with prior methods (CNNs, Transformers, and SSMs) on the Waymo Open Dataset validation set using 100% training data. Metrics are reported as mAP/mAPH (%) $\uparrow$ for overall results and AP/APH (%) $\uparrow$ per category. All models are trained under the single-frame setting. “–” indicates unavailable results. VDM-OD denotes a variant where the Only-Densification structure is used within the VDM framework. The best-performing results are highlighted in bold.
@@
CenterPoint [27] 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
Transfusion-LiDAR [1] 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
DSVT [24] 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.9 71.0
HEDNet [29] 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
SAFDNet [28] 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
LION [14] 72.1 68.0 87.9 64.9 77.6 44.4 28.5 89.6 75.6 59.4 80.8 71.6
FSHNet[13] 71.7 68.1 88.7 61.4 79.3 47.8 26.3 89.3 76.7 60.5 78.6 72.3
UniMamba [11] 72.6 68.5 88.7 64.7 79.7 47.9 28.7 89.7 74.6 59.1 79.5 72.3
@@
TransFusion-LiDAR [1] 70.2 65.5 86.2 56.7 66.3 58.8 28.2 86.1 68.3 44.2 82.0 78.2
DSVT [24] 72.7 68.4 86.8 58.4 67.3 63.1 37.1 88.0 73.0 47.2 84.9 78.4
HEDNet [29] 72.0 67.7 87.1 56.5 70.4 63.5 33.6 87.9 70.4 44.8 85.1 78.1
VPF [10] 72.7 67.0 85.8 55.1 63.5 62.1 33.3 87.6 72.5 48.6 82.9 78.2
SAFDNet [28] 72.3 68.3 87.3 57.3 68.0 63.7 37.3 89.0 71.1 44.8 84.9 79.5
LION [14] 73.9 69.8 87.2 61.1 68.9 65.0 36.3 90.0 74.0 49.2 87.3 79.5
UniMamba[11] 74.0 70.2 87.9 60.4 70.9 65.9 36.7 90.5 73.5 49.5 86.9 79.4
@@
VDM-OD (Ours) 73.9 70.5 88.3 62.2 70.7 65.3 35.0 90.6 74.4 51.1 88.3 78.8
TABLE III: Performances on the nuScenes validation and test set. ‘T.L.’, ‘C.V.’, ‘Ped.’, ‘M.T.’, ‘T.C.’, and ’B.R.’ are short for trailer, construction vehicle, pedestrian, motor, traffic cone, and barrier, respectively. VDM-OD denotes a variant where the Only-Densification structure is used within the VDM framework.
@@
CenterPoint [27] 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
HEDNet [29] 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
VoxelNeXt [5] 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
FSDV2 [7] 37.6 77.0 47.6 70.5 43.6 41.5 53.9 58.5 56.8 39.0 60.7 49.4 28.4 41.9 30.2 44.9 33.4 16.6 7.3 32.5 45.9 24.0 1.0 12.6 17.1 26.3 17.2
SAFDNet [28] 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
LION [14] 41.5 75.1 43.6 73.9 53.9 45.1 66.4 74.7 61.3 48.7 65.1 56.2 21.7 42.7 25.3 58.4 28.9 23.6 8.3 49.5 47.3 19.0 0.0 31.4 8.7 37.6 11.8
FSHNet[13] 40.2 77.4 48.3 72.9 50.0 47.2 63.4 69.9 56.1 43.8 62.5 53.7 31.8 44.8 28.9 56.6 32.0 23.0 8.7 34.5 41.0
UniMamba [11] 42.0 78.9 47.9 74.3 51.8 46.8 67.8 76.9 55.8 51.7 62.8 52.2 30.2 44.6 24.6 28.1 59.4 32.2 23.2 6.7 41.5 48.5 0.0 26.4 8.1 36.4 13.7
@@
TABLE IV: Comparison with prior methods on the Argoverse2 validation set. ‘Vehicle’, ‘C-Barrel’, ‘MPC-Sign’, ‘A-Bus’, ‘C-Cone’, ‘V-Trailer’, ‘MBT’, ‘W-Device’, and ‘W-Rider’ are abbreviations for regular vehicle, construction barrel, mobile pedestrian crossing sign, articulated bus, construction cone, vehicular trailer, message board trailer, wheeled device, and wheeled rider, respectively. VDM and VDM-OD refer to the proposed VDM-Mamba and VDM-OD variants.
IV Experiments
IV-A Datasets and Evaluation Metrics
Waymo Open Dataset. The Waymo Open Dataset (WOD) [21] includes 1,150 scenes, divided into 798 for training, 202 for validation, and 150 for testing. Each scene covers a perception range of $150\,\text{m}\times 150\,\text{m}$. For evaluation, WOD adopts 3D mean Average Precision (mAP) and mAP weighted by heading accuracy (mAPH). Each metric includes two difficulty levels: L1 for objects with more than five points, and L2 for bounding boxes containing one to five points.
nuScenes Dataset. The nuScenes dataset [3] contains 1,000 scenes, with 750 for training, 150 for validation, and 150 for testing. It uses mean Average Precision (mAP) and nuScenes Detection Score (NDS) as evaluation metrics. The dataset covers 10 annotated object classes, representing most common categories in traffic scenarios.
Argoverse 2 Dataset. The Argoverse 2 dataset [4] comprises 1,000 sequences, with 700 for training, 150 for validation, and 150 for testing. The perception range extends up to 200 meters. It uses mean Average Precision (mAP) as the evaluation metric.
ONCE Dataset. The ONCE dataset [16] consists of 5,000, 3,000, and 8,000 frames for training, validation, and testing respectively. It defines three detection classes: Vehicle, Pedestrian, and Cyclist. The final performance is evaluated using mean Average Precision (mAP) across all three classes.
IV-B Implementation Details
Network Architecture. We evaluate the proposed VDM module by integrating it with both the SSMs-based LION [14] and the Transformer-based DSVT [24] models. We use voxel sizes of $(0.075\,\text{m},0.075\,\text{m},0.25\,\text{m})$, $(0.08\,\text{m},0.08\,\text{m},0.1875\,\text{m})$, $(0.1\,\text{m},0.1\,\text{m},0.25\,\text{m})$, and $(0.1\,\text{m},0.1\,\text{m},0.25\,\text{m})$ for the nuScenes, Waymo, Argoverse 2, and ONCE datasets respectively. The number of voxel feature channels is set to 64. Specifically for the nuScenes dataset, we incorporate a Sparse Residual Module to enhance spatial feature interactions among voxels.
In the VDM module, the SubM3D (including SRB) layers use a kernel size of 3. All SparseConv3D layers adopt a kernel size of 3, stride of 2, and padding of 1 to downsample the feature map resolution. The channel configurations within VDM are set as $(64,32)$, $(32,64)$, and $(64,128)$. For LION and DSVT, we follow the default parameter settings from their original papers without modification. Additionally, we conduct an only-densification experiment, in which the voxel size along the $x$ and $y$ axes is scaled by a factor of 4, while the $z$ axis remains unchanged.
Training Process. We train all models using 8 NVIDIA A40 GPUs. On the Waymo Open Dataset, the VDM-SSMs configuration is trained for 24 epochs with a batch size of 16, while the VDM-DSVT configuration uses 24 epochs with a batch size of 8. For the remaining three datasets, we adopt the same training strategy as that used in the LION [14] baseline.
IV-C Main Results
@@
Results on WOD. As shown in Table II, we incorporate the VDM module into both Mamba-based and Transformer-based models. Among them, VDM-Mamba achieves a state-of-the-art (SOTA) result of 74.8 mAPH (L2), outperforming the corresponding baseline LION model (without VDM) by +0.8 mAPH, and also surpassing the recent UniMamba [11]. In addition, VDM-DSVT also demonstrates performance gains, achieving a +1.2 mAPH (L2) improvement over the original DSVT, which validates the general applicability of the VDM module to both SSM-based and Transformer-based detectors.
When only the densification mechanism is retained in VDM (i.e., VDM-OD), the model still achieves the best results on the Vehicle and Cyclist categories, highlighting the importance of the densification component in the overall design. Moreover, VDM-Mamba achieves the best performance in the Pedestrian category, suggesting that the fine-grained aggregation ability introduced by VDM is particularly beneficial for small object detection.
@@
PointRCNN [20] 52.1 74.5 40.9 16.8 4.3 6.2 2.4 0.9 29.8 46.0 20.9 5.5 28.7
PointPillars [12] 68.6 80.9 62.1 47.0 17.6 19.7 15.2 10.2 46.8 58.3 40.3 25.9 44.3
SECOND [25] 71.2 84.0 63.0 47.3 26.4 29.3 24.1 18.1 58.0 70.0 52.4 34.6 51.9
PV-RCNN [18] 77.8 89.4 72.6 58.6 23.5 25.6 22.8 17.3 59.4 71.7 52.6 36.2 53.6
CenterPoint [27] 66.8 80.1 59.6 43.4 49.9 56.2 42.6 26.3 63.5 74.3 57.9 41.5 60.1
PointPainting [23] 66.2 80.3 59.8 42.3 44.8 52.6 36.6 22.5 62.3 73.6 57.2 40.4 57.8
LION [14] 78.2 89.1 72.6 57.5 53.2 62.4 44.0 24.5 68.5 79.2 63.2 47.1 66.6
VDM-mamba (Ours) 78.6 89.1 74.3 59.3 54.2 63.3 43.2 26.1 69.9 79.4 64.9 49.8 67.6
VDM-OD (Ours) 78.5 88.8 73.1 59.7 51.0 59.2 42.3 24.5 68.7 79.1 64.1 48.9 66.1
TABLE V: Comparison with previous methods on ONCE validation set.
@@
Results on nuScenes. To further verify the effectiveness of the proposed VDM, we conduct experiments on both the nuScenes validation and test sets. As shown in Table III, on the validation set, VDM-Mamba achieves 68.1 mAP and 72.5 NDS, while the VDM-OD variant achieves even higher performance with 68.5 mAP and 72.9 NDS. Most notably, on the test set, VDM-OD achieves 70.5 mAP, surpassing the previous state-of-the-art UniMamba (70.2 mAP) and LION (69.8 mAP), while maintaining a competitive 73.9 NDS.
Detailed analysis reveals that VDM-Mamba excels in detecting vulnerable road users, achieving the best results in the Pedestrian and Bicycle categories on the validation set. This demonstrates that the fine-grained aggregation component in the full VDM offers clear advantages for small object localization. However, the superior overall performance of VDM-OD (70.5 mAP on Test) suggests that the densification-only structure plays the most critical role in mitigating feature sparsity and boosting general detection accuracy.
@@
Results on Argoverse V2. As shown in Table IV, our proposed models, VDM-Mamba and VDM-OD, achieve state-of-the-art detection performance on the Argoverse 2 validation set, reaching 42.3 mAP and 42.6 mAP, respectively. Compared with the strong baseline LION [14], VDM-Mamba yields a notable improvement of 0.8 mAP. Notably, VDM-OD, which retains only the densification structure within the VDM module, still surpasses all previous approaches, highlighting the effectiveness of the densification mechanism. These results further demonstrate the strong representational capacity and scalability of the VDM module in large-scale point cloud scenarios.
@@
TABLE VI: Ablation study on the ONCE and Waymo (20% data) dataset for fine-grained feature aggregation. Base denotes the detection results obtained by combining the VDM and Mamba network without fine voxel partitioning. Fine represents the results after applying fine-grained voxel partitioning.
Method Speed (s)
Waymo nuScenes Argoverse2 ONCE
LION 0.206 0.213 0.185 0.149
VDM 0.357 0.291 0.320 0.228
TABLE VII: Runtime comparison on different datasets. The inference speed is measured in seconds (s) on a single NVIDIA 3090 GPU with a batch size of 1.
Densification Waymo(3D AP/APH)L2
Vehicle Pedestrian Cyclist
No(LION) 68.8/68.3 77.3/72.6 74.7/73.8
Yes(VDM-OD) 70.3/69.9 78.2/73.8 76.0/75.0
TABLE VIII: Detection results of the voxel densification module on the Waymo validation set (20% data). “No” indicates that the original voxels are not diffused, and the detection results are entirely determined by LION [14]. “Yes” indicates that we only adopt the densification module from VDM, i.e., using sparse 3D spconv with a kernel size of 3 and a stride of 1.
Results on ONCE. VDM-Mamba demonstrates superior performance on the ONCE dataset (Table V). It achieves detection accuracies of 78.6 mAP, 54.2 mAP, and 69.9 mAP on the Vehicle, Pedestrian, and Cyclist categories, respectively. The overall detection accuracy reaches 67.6 mAP, representing a +1.0 mAP improvement over the baseline LION detector, establishing VDM-Mamba as the new state-of-the-art (SOTA).
In contrast to the nuScenes results, the VDM-OD variant (utilizing only densification) achieves 66.1 mAP, which is lower than the full VDM framework. This performance gap highlights the critical contribution of the fine-grained feature aggregation component on this dataset. While voxel expansion addresses sparsity, the aggregation module captures detailed local geometric cues that prove essential for the ONCE benchmarks, validating the necessity of our dual-path design (densification + aggregation) for robust detection across diverse scenarios.
Qualitative Visualization. To visually illustrate the advanced capabilities of our proposed model (VDM), we present a qualitative comparison of its detection performance against LION [14] on the Waymo validation set. As depicted in Figure 5, the annotations marked by yellow arrows highlight several advantages of our method. Firstly, the bounding boxes predicted by VDM exhibit a higher degree of alignment with the ground truth annotations. Secondly, VDM is capable of correctly detecting objects (true positives) that are entirely missed by LION. Finally, our model demonstrates a lower false positive rate, indicating superior precision.
Runtime Comparison. Table VII compares the inference time of our VDM with LION [14]. While VDM’s voxel densification approach increases latency by processing more voxels, it offers a favorable accuracy-efficiency trade-off. On the nuScenes and ONCE datasets, latency increases by only 0.08s for improvements of 0.4 in NDS and 1.0 in mAP. For long-range detection on datasets like Waymo and Argoverse 2, the more substantial latency is justified by significant performance gains of 0.8 mAPH (L2) and 0.8 mAP, respectively.
IV-D Ablation Study
Figure 5: A comparison of point cloud detection results between VDM and LION on the Waymo validation set. The blue, green, and red boxes represent human annotations, VDM predictions, and LION predictions, respectively. Yellow arrows on the figures highlight the differences in detection results. In the point cloud visualization process, the score threshold is set to 0.3, and the non-maximum suppression (NMS) threshold is set to 0.1.
Ablation Study on Voxel Densification. To assess the contribution of voxel densification to detection performance, we conduct ablation experiments on the Waymo validation set using a fixed voxel size of [0.32, 0.32, 0.1875]. We evaluate the model both with and without the densification mechanism. As shown in Table VIII, introducing voxel expansion improves the VDM-OD model by approximately 1.3 mAPH (L2), underscoring its critical role in sequential detection frameworks. Moreover, on the nuScenes (Table III) and Argoverse 2 (Table IV) datasets, we observe that under the Only-Densification setting, the model achieves impressive detection accuracy, further validating the significant impact of pre-serialization voxel expansion.